OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 12 -1

OpenStudy (zarkon):

replace x with 3

OpenStudy (anonymous):

so it would be 7?

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

f(3)= 3 +4, where x is replaced by 3, so its 7 yeah

OpenStudy (anonymous):

the same method applies to every function. For example if f(x) = x^3 +1, then f(1) = 1^3 +1 = 2

OpenStudy (anonymous):

thnx